OpenStudy (anonymous):

Hello, this is probability. Find the probability that in 10 tosses of a fair coin, there will be at least 2 heads. Can someone please help?

OpenStudy (anonymous):

P(at least 2 heads)=1-(P(1 head)+P(0 heads)) P(0 heads)=1/2^10 P(1 head)=10C1*1/2^10=10*1/2^10 thus, P(at least 2 heads)=1-11/(2^10)

OpenStudy (anonymous):

P(n heads)=10Cn*1/2^10

OpenStudy (anonymous):

oh okay. Thank you :)